OpenStudy (anonymous):

Check my work? -14x - 7(4 - x) <3(2x - 15) -14x - 28 - 7x < 6x - 45 -21x - 28 < 6x - 45 -21x + 21x - 28 < 6x + 21x - 45 -28 < 27x - 45 -28 + 45 < 27x - 45 + 45 17 < 27x 17/27 < 27/x 17/27 < x

OpenStudy (asnaseer):

you have made a mistake in the expansion of: - 7(4 - x) remember that two minuses make a plus.

OpenStudy (anonymous):

-number multiply by a negative number makes positive @karldwalsh

OpenStudy (anonymous):

- 7(4 - x) = -28 +7x

OpenStudy (asnaseer):

a(b - c) = ab - ac -a(b - c) = -ab + ac [because (-a) times (-c) equals ac]